OpenStudy (anonymous):

write the explicit definition for the sequence -2, 1, 4, 7... a. -2+3n b.-2+(n-3) c.-2+(n+3) d.-2+3(n-1)

OpenStudy (trojanpoem):

a + c(n-1) c = 1 - (-2) = 3 -2 +3(n-1)